OpenStudy (anonymous):

What is the solution of the system? Use substitution. 3x + 2y = 7 –3x + 11 = y

OpenStudy (anonymous):

If you just want answers you can simply type both equations into wolfram alpha (www.wolframalpha.com) and it gives you the answer

OpenStudy (anonymous):

2y + 11 = y + 7 y = -4

OpenStudy (anonymous):

But it's x=5, y=-4